### v1

@@ Abstract
This work introduces efficient and accurate spectral solvers for nonlocal equations on bounded domains. These spectral solvers exploit the fact that integration in the nonlocal formulation transforms into multiplication in Fourier space and that nonlocality is decoupled from the grid size. As a result, the computational cost is reduced to $\mathcal{O}(N\log N)$ for an $N$-point discretization grid. Our approach extends the spectral solvers developed in [1] for periodic domains by incorporating the two-dimensional Fourier continuation (2D-FC) algorithm introduced in [7]. We evaluate the performance of the proposed methods on two-dimensional nonlocal Poisson and nonlocal diffusion equations defined on bounded domains. While the regularity of solutions to these equations in bounded settings remains an open problem, we conduct numerical experiments to explore this issue, particularly focusing on studying discontinuities.
@@ 1 Introduction and overview
[…]
@@
Nonlocal operators of this form have their roots in peridynamics [5, 33, 34, 35] and they have been used in different applied settings, such as fractures [20, 6, 12, 26], image processing [18], tumor growth [25], material failure and damage [22, 13, 21], among others. There have been several mathematical and computational results involving nonlocal equations, see for example [11, 16, 17, 19, 28, 29, 9, 27]. On the mathematical side, the work in [2] introduces the Fourier multipliers for nonlocal Laplace operators, studies the asymptotic behavior of these multipliers, and then applies the asymptotic analysis in the periodic setting to prove regularity results for the nonlocal Poisson equation. The Fourier multipliers $m^{\delta,\beta}$ of the nonlocal operator $\mathcal{L}^{\delta,\beta}$ are defined through Fourier transform by
@@
These multipliers are given by the integral representation (2), and they are used to obtain the eigenvalues of the nonlocal operator $\mathcal{L}^{\delta,\beta}$ in the case of periodic domains by evaluating them on a lattice of points (see Section 4.1 of [2]).
[…]
As pointed out in [14, 15], it is challenging to accurately and efficiently compute the eigenvalues of nonlocal operators. The work in [1] simplifies the computation of these multipliers by representing the $n$-dimensional integral in (2) by a hypergeometric function given by
@@
The work in [30] and [10] apply these Fourier multipliers to study the regularity of solutions for nonlocal diffusion equations and nonlocal wave equations over the space of periodic distributions, respectively. As emphasized in [30], nonlocal diffusion equation satisfies an instantaneous smoothing effect when the integral kernel is singular with $\beta>n$ and a gradual (over time) smoothing effect when $\beta=n$. However, for integrable kernels ($\beta<n$), the nonlocal diffusion equation is non-smoothing. According to the work in [2], the regularity of the solution $u$ for the nonlocal Poisson equation $\mathcal{L}^{\delta,\beta}u=f$ is the same as the regularity of $f$ when $\beta\leq n$, while when $\beta>n$, the solution is $\beta-n$ more regular than $f$. Specifically, if $f\in H^{s}(T^{n})$ for $s\geq 0$, then the solution $u\in H^{s}(T^{n})$ when $\beta\leq n$ and $u\in H^{s+\beta-n}(T^{n})$ when $\beta>n$.
@@
For computational results, on the other hand, many computational techniques and numerical analysis have been developed for solving nonlocal equations, see for example [9, 11, 15, 3, 8, 14, 1, 36]. Spectral methods are accurate and efficient for solving nonlocal equations because they transform the integral in the nonlocal formulation to multiplication in Fourier space. For example, the application of the operator $\mathcal{L}^{\delta,\beta}$ on a function $u$ results in multiplication in Fourier space, as shown in (1), and this runs at a computational cost of $\mathcal{O}(N\log N)$ for an $N$-point discretization grid. For comparison of the efficiency and accuracy of spectral methods versus standard numerical methods such as finite differences, see [1].
@@
Some spectral methods for nonlocal equations have been developed in [14, 1, 23, 15, 36]. The work in [23] introduces a boundary-adapted spectral method for peridynamic diffusion problems with arbitrary boundary conditions. A spectral method for nonlocal equations on periodic domains has also been developed in [1] with high accuracy. One difficulty that arises in adapting the spectral method developed in [1] to solve nonlocal equations on bounded domains is that the solutions are in general nonperiodic, therefore exhibiting the Gibbs phenomenon as the Fourier coefficients of the solution decay slowly. One way around this is to use Fourier continuation. The idea behind Fourier continuation is to reduce the impact of slowly decaying Fourier coefficients of a smooth but nonperiodic function on a given domain by constructing a smooth periodic extension of the function over a larger domain.
@@
Our proposed solvers build on the Fourier spectral method developed in [1] for periodic domains and incorporate the two-dimensional Fourier continuation (2D-FC) algorithm from [7] to solve two-dimensional nonlocal equations on bounded domains. We also use this approach to perform numerical experiments aimed at exploring the regularity of solutions for nonlocal equations, with particular focus on discontinuities. For integrable kernels, numerical simulations indicate that discontinuities may develop in the solution. In contrast, for singular kernels with $\beta>n+1=3$, the solution appear to remain continuous in the interior of the domain, and no discontinuities are observed.
@@
The paper is organized as follows. In Section 2, we describe the two-dimensional Fourier continuation (2D-FC) algorithm developed in [7], along with minor modifications. Section 3 presents our spectral solvers for two-dimensional nonlocal Poisson and nonlocal diffusion equations on bounded domains. In Section 4, we discuss the numerical results and explore the behavior of solutions in various regimes.
@@ 2 Two dimensional Fourier continuation
In this section, we provide a description of the 2D-FC algorithm presented in [7]. The description includes a modification of some of the steps in the algorithm to improve efficiency and accuracy. In particular, we modify the steps in computing the continuation values on the Cartesian grids in the extended region. For consistency, we follow a notation similar to that of [7].
@@ 2.1 Overview of 1D blending-to-zero algorithm
In the description given in [4], the blending-to-zero algorithm starts by smoothly extending the vector $f=(f_{1},f_{2},\dots,f_{N})\in\mathbb{R}^{N}$ of values of a complex-valued function $f$ over the set of uniformly spaced nodes $\mathcal{D}=\{x_{1},x_{2},\dots,x_{N}\}$ where for a given step size $h>0$, $x_{j}=x_{1}+(j-1)h.$ Without loss of generality, we assume that $x_{1}=0$ and $x_{N}=1$. The extension relies on some number $d$ of near-boundary function values, $\{f_{1},f_{2},\dots,f_{d}\}$ near $x=0$ and $\{f_{N-d+1},f_{N-d+2},\dots,f_{N}\}$ near $x=1$, and it produces a sequence $\{\tilde{f}_{j}\}_{j=-\infty}^{\infty}$ with the property that
@@
The extended vector $\tilde{f}$ is then blended to zero on either end by multiplying with a suitable windowing function. A complete description of this algorithm is provided in [4]. In Figure 1, we demonstrate the 1D blending to zero of the nonperiodic function $f:[0,1]\to\mathbb{R}$ given by $f(x)=\sin(15x)e^{x}$.
Figure 1: A demonstration of the 1D blending to zero on a smooth but nonperiodic function $f(x)=\sin(15x)e^{x}$ defined on [0,1]. The black solid points denote the $d$ near-boundary function values used to produce the extension of $f$ denoted by the blue solid points.
@@
$f^{c}(x,y)=\sum_{k=\frac{N_{x}}{2}+1}^{N_{x}/2}\sum_{l=\frac{N_{y}}{2}+1}^{N_{y}/2}\hat{f}^{c}_{kl}\exp\left(\frac{2\pi ikx}{L_{x}}\right)\exp\left(\frac{2\pi ily}{L_{y}}\right),$ (4)
@@ 2.2.1 Curvilinear grid along a 2D tangential-normal
where $q(\theta)=(x(\theta),y(\theta))$ is the parametrization of $\Gamma$, $n(\theta)=(n_{x}(\theta),n_{y}(\theta))$ is the unit normal to the boundary at $q(\theta)$, and $d,C,k_{1}$ represent the number of matching points near the boundary, the number of continuation points, and the step size along the normals respectively. Consider the following uniform discretization of the interval $[0,2\pi]$
@@ 2.2.2 Computing continuation values on $V^{+}_{\theta_{p}}$
For a given $\theta_{p}\in\Theta_{B}$, a continuation of $f$ on $V^{+}_{\theta_{p}}$ is obtained via the blending-to-zero algorithm described in Section 2.1. We first obtain the approximations of $f$ on the set $V^{-}_{\theta_{p}}$, which are indicated by black solid circles in Figure 3, and then use these approximations to obtain a continuation of $f$ via the 1D blending-to-zero procedure along the corresponding normal direction. The values of $f$ on $V^{-}_{\theta_{p}}$ are obtained by a two-step polynomial interpolation scheme, using polynomials of degree $M-1$. This is briefly described as follows:
Consider the case where $|n_{x}(\theta_{p})|\geq|n_{y}(\theta_{p})|$. We first estimate the values of $f$ at each red point, which is the intersection of the normal line and the vertical Cartesian grid lines as shown in Figure 3, by interpolating vertically the function values at the selected $M$ Cartesian points, shown as open circles, onto the red points. Then we interpolate the values of $f$ at the $M$ red points onto the black solid circles, using a polynomial of degree $M-1$.
@@ 2.2.3 Computing continuation values on the Cartesian grid
Once all the continuation values have been obtained on $V_{\Theta_{B}}^{+}=\{V_{\theta_{p}}^{+},\theta_{p}\in\Theta_{B}\}$, we then obtain the continuation of $f$ on the Cartesian grid points within $V^{+}$ using a two-step 1D polynomial interpolation scheme. The 2D-FC algorithm is then completed by evaluating corresponding FCs in (4) by means of 2D FFT. The description of the two-step 1D polynomial interpolation scheme is presented below.
@@
where $f^{c}(b_{j})$ is the continuation of $f$ at $b_{j}$ obtained via a 1D polynomial interpolation on $M$ blending to zero points along $n(\theta_{j})$ (see Figure 4(b)). Clearly, $\mathcal{T}(0)=f^{c}(Q)$. It then follows that a 1D polynomial interpolation procedure on the function $\mathcal{T}(\tau)$, using polynomial of degree $M-1$, can be used to obtain the desired continuation of $f$ at Cartesian point $Q$.
@@ 2.2.4 Parameter description and selection
  • $d$: number of near-boundary points used to blend $f$ to zero along each normal (subsection 2.1).
  • $C$: number of continuation points in the 1D blending to zero process (subsection 2.1).
  • $k_{1}$: normal step size (subsection 2.2.1)
  • $k_{2}$: boundary step size (subsection 2.2.1)
  • $h$: step size used to construct Cartesian grid $\mathcal{H}$ (subsection 2.2.1)
@@
  • $N=N_{x}\times N_{y}$: number of points in the uniform Cartesian grid (subsection 2.2.1)
  • $B$: number of boundary discretization points (subsection 2.2.1)
  • $M-1$: the degree of the interpolating polynomial (subsections 2.2.2 and 2.2.3)
For a given step size $h$ used to construct the Cartesian grid $\mathcal{H}$, we choose normal step size $k_{1}$ and boundary step size $k_{2}$ to coincide with $h$. The parameters $C=25$ and $n_{r}=6$ were used in the 1D blending to zero procedure described in Subsection 2.1. Finally, we set the interpolating polynomial degree $(M-1)=d$ for all matching point numbers $d$ considered except the interpolation-degree experiments presented in Example 1.
In the following example, we demonstrate the performance of the 2D-FC method using three interpolation degrees: $M-1=d,$ $d+1$, and $d+2$ for a given matching point number $d$.
@@
We consider the FC approximation of the function $f:\Omega\to\mathbb{R}$ given by
@@
on the unit disk $\Omega=\{(x,y):x^{2}+y^{2}\leq 1\}$ with three different values of $d$: $d=4$, $5$, and $8$; and three different values of $M$: $M=d+1$, $d+2$, and $d+3$.
Figure 5 displays the relative $L^{2}$ errors as functions of the number of spatial discretization, $N_{1D}$ ($N_{x}=N_{y}\approx N_{1D}+2C$), in one direction within $\Omega$. The errors are computed on a Cartesian grid of step size $h/2$ within $\Omega$. The approximations of $f$ on the finer grid are obtained by means of zero padding [31]. For the case $M=d+1$ (Figure 5(a)), we observe fifth, sixth, and ninth orders of convergence for $d=4$, $d=5$, and $d=8$ respectively, showing convergence of order $d+1$. However, for the case $M=d+2$ (Figure 5(b)), we observe order $d+2$ convergence for $d=5$ and $d=8$, and order $d+1$ convergence for $d=4$. For the case $M=d+3$ (Figure 5(c)), we observe order $d+1$ convergence for $d=4$ and $d=5$, and order $d+2$ convergence for $d=8$. The case $M=d+1$, which gives a similar order of convergence ($d+1$) for the three values of $d$ considered, will be used in all of the numerical experiments presented in this work.
@@
Figure 5: Convergence study for the 2D-FC approximation of the function $f$ in (6) with three choices of matching point numbers $d$: $d=4$, $5$, and $8$; and three choices of $M$: $M=d+1$, $d+2$, and $d+3$. The integer $N_{1D}$ is the number of spatial grid points in one spatial direction within $\Omega$.
@@
over the domain defined in (7). Figure 6 displays both the original function and its extension.

### v2

@@ Abstract
This work introduces efficient and accurate spectral solvers for nonlocal equations on bounded domains. These spectral solvers exploit the fact that integration in the nonlocal formulation transforms into multiplication in Fourier space and that nonlocality is decoupled from the grid size, allowing fast and accurate solutions to the nonlocal problems. Our approach extends the spectral solvers developed in [alali2020fourier] for periodic domains by incorporating the two-dimensional Fourier continuation (2D-FC) algorithm introduced in [bruno2022two]. We evaluate the performance of the proposed methods on two-dimensional nonlocal Poisson and nonlocal diffusion equations defined on bounded domains. While the regularity of solutions to these equations in bounded settings remains an open problem, we conduct numerical experiments to explore this issue, particularly focusing on studying discontinuities.
@@ 1 Introduction and overview
$\mathcal{L}^{\delta,\beta}u(x)=c^{\delta,\beta}\int_{B_{\delta}(x)}\frac{u(y)-u(x)}{\|y-x\|^{\beta}}dy,~x\in\mathbb{R}^{n},$
@@
Nonlocal operators of this form have their roots in peridynamics [bobaru2016handbook, silling2000reformulation, silling2007peridynamic, silling2017modeling] and they have been used in different applied settings, such as fractures [hu2012peridynamic, bobaru2015cracks, dias2017review, lipton2014dynamic], image processing [gilboa2009nonlocal], tumor growth [lejeune2017modeling], material failure and damage [isiet2021review, diehl2015simulation, hu2017peridynamics], among others. There have been several mathematical and computational results involving nonlocal equations, see for example [de2017finite, foss2016differentiability, foss2018existence, hinds2012dirichlet, mengesha2014nonlocal, mengesha2016characterization, burkovska2024efficient, madenci2014coupling]. On the mathematical side, the work in [alali2021fourier] introduces the Fourier multipliers for nonlocal Laplace operators, studies the asymptotic behavior of these multipliers, and then applies the asymptotic analysis in the periodic setting to prove regularity results for the nonlocal Poisson equation. The Fourier multipliers $m^{\delta,\beta}$ of the nonlocal operator $\mathcal{L}^{\delta,\beta}$ are defined through Fourier transform by
@@
These multipliers are given by the integral representation (2), and they are used to obtain the eigenvalues of the nonlocal operator $\mathcal{L}^{\delta,\beta}$ in the case of periodic domains by evaluating them on a lattice of points (see Section 4.1 of [alali2021fourier]).
$m^{\delta,\beta}(\nu)=c^{\delta,\beta}\int_{B_{\delta}(0)}\frac{\cos(\nu\cdot z)-1}{\|z\|^{\beta}}dz.$ (2)
As pointed out in [du2016asymptotically, du2017fast], it is challenging to accurately and efficiently compute the eigenvalues of nonlocal operators. The work in [alali2020fourier] simplifies the computation of these multipliers by representing the $n$-dimensional integral in (2) by a hypergeometric function given by
@@
The work in [mustapha2023regularity] and [dang2024regularity] apply these Fourier multipliers to study the regularity of solutions for nonlocal diffusion equations and nonlocal wave equations over the space of periodic distributions, respectively. As emphasized in [mustapha2023regularity], nonlocal diffusion equation satisfies an instantaneous smoothing effect when the integral kernel is singular with $\beta>n$ and a gradual (over time) smoothing effect when $\beta=n$. However, for integrable kernels ($\beta<n$), the nonlocal diffusion equation is non-smoothing. According to the work in [alali2021fourier], the regularity of the solution $u$ for the nonlocal Poisson equation $\mathcal{L}^{\delta,\beta}u=f$ is the same as the regularity of $f$ when $\beta\leq n$, while when $\beta>n$, the solution is $\beta-n$ more regular than $f$. Specifically, if $f\in H^{s}(T^{n})$ for $s\geq 0$, then the solution $u\in H^{s}(T^{n})$ when $\beta\leq n$ and $u\in H^{s+\beta-n}(T^{n})$ when $\beta>n$.
@@
For computational results, on the other hand, many computational techniques and numerical analysis have been developed for solving nonlocal equations, see for example  [burkovska2024efficient, de2017finite, du2017fast, albin2011spectral, burkovska2020affine, du2016asymptotically, alali2020fourier, slevinsky2018spectral]. Spectral methods are accurate and efficient for solving nonlocal equations because they transform the integral in the nonlocal formulation to multiplication in Fourier space. For example, the application of the operator $\mathcal{L}^{\delta,\beta}$ on a function $u$ results in multiplication in Fourier space, as shown in (1), and this runs at a computational cost of $\mathcal{O}(N\log N)$ for an $N$-point discretization grid. For comparison of the efficiency and accuracy of spectral methods versus standard numerical methods such as finite differences, see [alali2020fourier].
@@
Some spectral methods for nonlocal equations have been developed in [du2016asymptotically, alali2020fourier, jafarzadeh2020efficient, du2017fast, slevinsky2018spectral]. The work in [jafarzadeh2020efficient] introduces a boundary-adapted spectral method for peridynamic diffusion problems with arbitrary boundary conditions. A spectral method for nonlocal equations on periodic domains has also been developed in [alali2020fourier] with high accuracy. One difficulty that arises in adapting the spectral method developed in [alali2020fourier] to solve nonlocal equations on bounded domains is that the solutions are in general nonperiodic, therefore exhibiting the Gibbs phenomenon as the Fourier coefficients of the solution decay slowly. One way around this is to use Fourier continuation. The idea behind Fourier continuation is to reduce the impact of slowly decaying Fourier coefficients of a smooth but nonperiodic function on a given domain by constructing a smooth periodic extension of the function over a larger domain.
@@
Our proposed solvers build on the Fourier spectral method developed in [alali2020fourier] for periodic domains and incorporate the two-dimensional Fourier continuation (2D-FC) algorithm from [bruno2022two] to solve two-dimensional nonlocal equations on bounded domains. We also use this approach to perform numerical experiments aimed at exploring the regularity of solutions for nonlocal equations, with particular focus on discontinuities. For integrable kernels, numerical simulations indicate that discontinuities may develop in the solution. In contrast, for singular kernels with $\beta>n+1=3$, the solution appear to remain continuous in the interior of the domain, and no discontinuities are observed.
@@
The paper is organized as follows. In Section 2, we describe the two-dimensional Fourier continuation (2D-FC) algorithm developed in [bruno2022two], along with minor modifications. Section 3 presents our spectral solvers for two-dimensional nonlocal Poisson and nonlocal diffusion equations on bounded domains. In Section 4, we discuss the numerical results and explore the behavior of solutions in various regimes. In Section 5, we provide a summary of our approach and outline directions for future research.
@@ 2 Two dimensional Fourier continuation
In this section, we provide a description of the 2D-FC algorithm presented in [bruno2022two]. The description includes a modification of some of the steps in the algorithm to improve efficiency and accuracy. In particular, we modify the steps in computing the continuation values on the Cartesian grids in the extended region. For consistency, we follow a notation similar to that of [bruno2022two].
@@ 2.1 Overview of 1D blending-to-zero algorithm
In the description given in [albin2014discrete], the blending-to-zero algorithm starts by smoothly extending the vector $f=(f_{1},f_{2},\dots,f_{N})\in\mathbb{R}^{N}$ of values of a complex-valued function $f$ over the set of uniformly spaced nodes $\mathcal{D}=\{x_{1},x_{2},\dots,x_{N}\}$ where for a given step size $h>0$, $x_{j}=x_{1}+(j-1)h.$ Without loss of generality, we assume that $x_{1}=0$ and $x_{N}=1$. The extension relies on some number $d$ of near-boundary function values, $\{f_{1},f_{2},\dots,f_{d}\}$ near $x=0$ and $\{f_{N-d+1},f_{N-d+2},\dots,f_{N}\}$ near $x=1$, and it produces a sequence $\{\tilde{f}_{j}\}_{j=-\infty}^{\infty}$ with the property that
@@
The extended vector $\tilde{f}$ is then blended to zero on either end by multiplying with a suitable windowing function. A complete description of this algorithm is provided in [albin2014discrete]. In Figure 1, we demonstrate the 1D blending to zero of the nonperiodic function $f:[0,1]\to\mathbb{R}$ given by $f(x)=\sin(15x)e^{x}$.
Figure 1: A demonstration of the 1D blending to zero on a smooth but nonperiodic function $f(x)=\sin(15x)\exp(x)$ defined on [0,1]. The black solid points denote the $d$ near-boundary function values used to produce the extension of $f$ denoted by the blue solid points.
@@
$f^{c}(x,y)=\sum_{k=-\frac{N_{x}}{2}+1}^{N_{x}/2}\sum_{l=-\frac{N_{y}}{2}+1}^{N_{y}/2}\hat{f}^{c}_{kl}\exp\left(\frac{2\pi ikx}{L_{x}}\right)\exp\left(\frac{2\pi ily}{L_{y}}\right),$ (4)
@@ 2.2.1 Curvilinear grid along a 2D tangential-normal
where $q(\theta)=(x(\theta),y(\theta))$ is the parametrization of $\Gamma$, $n(\theta)=(n_{x}(\theta),n_{y}(\theta))$ is the unit normal to the boundary at $q(\theta)$, and $d,C,k_{1}$ represent the number of matching points near the boundary, the number of continuation points, and the step size along the normals respectively. Consider the following uniform discretization of the interval $0,2\pi]$
@@ 2.2.2 Computing continuation values on $V^{+}_{\theta_{p}}$
For a given $\theta_{p}\in\Theta_{B}$, a continuation of $f$ on $V^{+}_{\theta_{p}}$ is obtained via the blending-to-zero algorithm described in Section 2.1. We first obtain the approximations of $f$ on the set $V^{-}_{\theta_{p}}$, which are indicated by black solid circles in Figure 3, and then use these approximations to obtain a continuation of $f$ via the 1D blending-to-zero procedure along the corresponding normal direction. The values of $f$ on $V^{-}_{\theta_{p}}$ are obtained by a two-step polynomial interpolation scheme, using polynomials of degree M1M-1. This is briefly described as follows:
Consider the case where $|n_{x}(\theta_{p})|\geq|n_{y}(\theta_{p})|$. We first estimate the values of $f$ at each red point, which is the intersection of the normal line and the vertical Cartesian grid lines as shown in Figure 3, by interpolating vertically the function values at the selected $M$ Cartesian points, shown as open circles, onto the red points. Then we interpolate the values of $f$ at the $M$ red points onto the black solid circles, using a polynomial of degree M1M-1.
@@ 2.2.3 Computing continuation values on the Cartesian grid
Once all the continuation values have been obtained on $V_{\Theta_{B}}^{+}=\{V_{\theta_{p}}^{+}:\theta_{p}\in\Theta_{B}\}$, we then obtain the continuation of $f$ on the Cartesian grid points within $V^{+}$ using a two-step 1D polynomial interpolation scheme. The 2D-FC algorithm is then completed by evaluating corresponding FCs in (4) by means of 2D FFT. The description of the two-step 1D polynomial interpolation scheme is presented below.
@@
where $f^{c}(b_{j})$ is the continuation of $f$ at $b_{j}$ obtained via a 1D polynomial interpolation on $M$ blending to zero points along $n(\theta_{j})$ (see Figure 4(b)). Clearly, $\mathcal{T}(0)=f^{c}(Q)$. It then follows that a 1D polynomial interpolation procedure on the function $\mathcal{T}(\tau)$, using polynomial of degree M1M-1, can be used to obtain the desired continuation of $f$ at Cartesian point $Q$.
@@
Remark 2.
We observe that when the value of $M$ becomes too large, the accuracy decreases. This loss of accuracy arises because using a larger number of normals in the interpolation procedure for estimating the point $Q$ (see Figure 4) causes the corresponding blue points to move closer to the boundary of $\Omega$. Consequently, the estimation of $Q$ is affected, since the values of the blue points near the boundary are not sufficiently blended to zero. To avoid this, we select a smaller number of normals (corresponding to a lower $M$) so that the contributing normals remain close to the point $Q$.
@@ 2.2.4 Parameter description and selection
  • $d$: number of near-boundary points used to blend $f$ to zero along each normal.
  • $C$: number of continuation points in the 1D blending to zero process.
  • $k_{1}$: normal step size.
  • $k_{2}$: boundary step size.
  • $h$: step size used to construct Cartesian grid $\mathcal{H}$.
@@
  • $N=N_{x}\times N_{y}$: number of points in the uniform Cartesian grid
  • $B$: number of boundary discretization points.
  • M1M-1: the degree of the interpolating polynomial.
For a given step size $h$ used to construct the Cartesian grid $\mathcal{H}$, we choose normal step size $k_{1}$ and boundary step size $k_{2}$ to coincide with $h$. The parameters $C=25$ and $n_{r}=6$ were used in the 1D blending to zero procedure described in Subsection 2.1. Finally, we set the interpolating polynomial degree (M1)=d(M-1)=d for all matching point numbers $d$ considered except the interpolation-degree experiments presented in Example 1.
In the following example, we demonstrate the performance of the 2D-FC method using three interpolation degrees: M1=d,M-1=d, $d+1$, and $d+2$ for a given matching point number $d$.
@@
We consider the FC approximation of a function $f:\Omega\to\mathbb{R}$ given by
@@
on the unit disk $\Omega=\{(x,y):x^{2}+y^{2}\leq 1\}$ with three different values of $d$: $d=4$, 55, and $8$; and three different values of $M$: $M=d+1$, $d+2$, and $d+3$.
Figure 5 displays the relative $L^{2}$ errors $\mathcal{E}_{L^{2}}^{\text{rel}}$, given in (7), as functions of the number of spatial discretization, $N_{1D}$ ($N_{x}=N_{y}\approx N_{1D}+2C$), in one direction within the domain $\Omega=\{(x,y):x^{2}+y^{2}\leq 1\}$. The errors are computed on a Cartesian grid of step size $h/2$ within $\Omega$. The approximations of $f$ on the finer grid are obtained by means of zero padding [press2007numerical]. Specifically, we compute $\mathcal{E}_{L^{2}}^{\text{rel}}$ using the following formula:
$\mathcal{E}_{L^{2}}^{\text{rel}}=\frac{\|f-f^{\mathrm{zp}}\|_{L^{2}(\Omega)}}{\|f\|_{L^{2}(\Omega)}},$ (7)
where $f^{\mathrm{zp}}$ denotes the zero–padded Fourier approximation of $f$, obtained by extending $f$ through Fourier continuation, then applying zero padding in Fourier space, and restricting the result back to $\Omega$.
For the case $M=d+1$ (Figure 5(a)), we observe convergence of orders five, six, and nine for $d=4$, $d=5$, and $d=8$ respectively, corresponding to an overall convergence rate of $\mathcal{O}(h^{d+1})$. However, for the case $M=d+2$ (Figure 5(b)), the observed rates are $\mathcal{O}(h^{d+1})$ for $d=4$ and $\mathcal{O}(h^{d+2})$ for $d=5$ and $d=8$. Similarly, for $M=d+3$ (Figure 5(c)), we observe $\mathcal{O}(h^{d+1})$ convergence for $d=4$ and $d=5$ and $\mathcal{O}(h^{d+2})$ convergence for $d=8$. In particular, we observe a lower convergence rate for higher values of $M$ as explained in Remark 2. The choice $M=d+1$, which yields a consistent convergence order of $d+1$ for the three values of $d$ considered, will be employed in all subsequent numerical experiments.
@@
Figure 5: Convergence study for the 2D-FC approximation of the function $f$ in (6) with three choices of matching point numbers $d$: $d=4$, 55, and $8$; and three choices of $M$: $M=d+1$, $d+2$, and $d+3$. The integer $N_{1D}$ is the number of spatial grid points in one spatial direction within $\Omega$.
@@
over the domain defined in (8).
Figure 6 shows the original function together with its periodic extension obtained using the 2D-FC method. The corresponding relative $L^{2}$ error $\mathcal{E}_{L^{2}}^{\text{rel}}$, computed using the formula in (7), is $1.1674\cdot 10^{-06}$.